# Zero Inflation as a Missing Data Problem: a Proxy-based Approach

### v1

@@ 3.3 Partial Identification in ZI MCAR
Consider a ZI MCAR model in Fig. 2 (a) under proxy assumptions A1, A2, with categorical $X$ and binary $R,W$. Given a consistent observed law $p(X,W)$ satisfying positivity assumption, the set of compatible proxy-indicator conditionals $q(W\mid R)$ is given by
@@ 3.4 Partial Identification in ZI MAR
Consider a ZI MAR model in Fig. 2 (b) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity, the set of compatible proxy-indicator conditional distributions $q(W\mid R,C)$ is given by, for each value $c$,
@@
Consider a ZI MAR model in Fig. 2 (c) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity, the set of compatible proxy-indicator conditional distributions $q(W\mid R)$ is given by
@@ 4 Experiments
We confirmed the validity of our analytical results for inflated zero models by sampling data generating processes (DGPs), and numerical methods. In addition, we used our methods to perform sensitivity analyses on CLABSI data. Details of these experiments are in the Appendix.
@@ 4.3 Data Application
Our working model is the proxy-augmented ZI MAR under assumptions A1 and A2, shown in Fig. 2 (b). Using the analytic bounds for the ZI MAR model derived in Section 3.4, we perform a sensitivity analysis to understand how the true CLABSI rate $p(X^{(1)}=1)$ changes as the proxy-indicator relationship $p(W\mid R)$ varies, within its compatibility range. First, we use the EM algorithm [Dempster et al., 1977] to maximize the observed data likelihood ^$\hat{\cal L}_{EM}(X,W,C)$ defined via the full data distribution consistent with our assumptions. Next, we invoke Theorem 5 to obtain the plug-in estimate for $p(W=0|R=1)$ and the bounds for $p(W=0|R=0)$. Finally, we do a grid search over the bounds interval, compute the full data distribution $p(R,X,W,C)$ for each value of $p(W=0|R=0)$ via (2) and obtain $p(X^{(1)})$ using standard g-formula adjustment in MAR models. The sensitivity analysis curve is shown in Fig 4.
The values of $p(W=0|R=0)$ consistent with the model show that inability to make a CLABSI determination is strongly associated with access to patient data via electronic health records. Our obtained (sharp) bound for the nuisance parameter $p(W=0|R=0)$ is $[0.89,1]$, yielding the estimated range of the true CLABSI rate to be $[0.68,0.78]$. Compared with the baseline rate of $65\%$ under no-zero-inflation assumption, the rate’s bound implies that anywhere from $3\%$ to $13\%$ of true CLABSI cases are undercounted, even in our patient cohort with a highly elevated CLABSI prevalence.
We have repeated the analysis using the proxy-augmented ZI MAR model under assumptions A1 and A2, shown in Fig. 2 (c). In this case, bounds for $p(W=0|R=0,c)$ were obtained, for each value $c$. The narrowest bound $[0.9,1.0]$ corresponds to patients receiving outpatient parenteral antimicrobial therapy via a peripherally inserted central catheter. On the other hand, the widest bound $[0.67,1.0]$ corresponds to patients receiving chemotherapy via PICC, a particular type of catheter. We performed a search over the polytope comprised of these bounds and find the estimated range of true CLABSI rate to be $[0.68,1.0]$. That is, anywhere from 3%-35% of true CLABSI cases are undercounted.
All bounds derived by our methods for the true CLABSI rate were deemed to be medically plausible by our medical collaborators.
@@ Acknowledgements.
This research is funded in part by ONR N00014-21-1-2820, NSF 2040804, NSF CAREER 1942239, NIH R01 AI127271-01A1, NIH R01 HS027819.
@@ A.4.2 Bounds for ZI MCAR
Due to this lemma, an observed law $p(X,W)$ is consistent to the model if and only if $\forall x\neq 0,p(W=0,X=x)=p(W=0,X=1)$. We also require positivity $\forall x,p(X=x)>0$.
Theorem 3 Consider a ZI MCAR model in Fig. 2 (a) (reproduced in Fig. 5) under proxy assumptions A1, A2, with categorical $X$ and binary $R,W$. Given a consistent observed law $p(X,W)$ satisfying positivity, the set of compatible proxy-indicator conditionals $q(W\mid R)$ is given by
@@ A.5 ZI MAR proofs
Theorem 4 Consider a ZI MAR model in Fig. 2 (b) (reproduced in Fig. 6) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity, the set of compatible proxy-indicator conditional distributions $q(W\mid R,C)$ is given by, for each value $c$,
@@
$q\left(X^{(1)},R,X,W,C\right)=q\left(X^{(1)},R,X,W\mid C\right)p(C)$ (43)
@@
Theorem 5 Consider a ZI MAR model in Fig. 2 (c) (reproduced in Fig. 7) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity, the set of compatible proxy-indicator conditional distributions $q(W\mid R)$ is given by

### v2

@@ 3.3 Partial Identification in ZI MCAR
Consider a ZI MCAR model in Fig. 2 (a) under proxy assumptions A1, A2, with categorical $X$ and binary $R,W$. Given a consistent observed law $p(X,W)$ satisfying positivity assumption, $\forall x,p(x)>0$, the set of compatible proxy-indicator conditionals $q(W\mid R)$ is given by
@@ 3.4 Partial Identification in ZI MAR
Consider a ZI MAR model in Fig. 2 (b) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity assumption, $\forall x,c,p(x,c)>0$, the set of compatible proxy-indicator conditional distributions $q(W\mid R,C)$ is given by, for each value $c$,
@@
Consider a ZI MAR model in Fig. 2 (c) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity assumption, $\forall x,c,p(x,c)>0$, the set of compatible proxy-indicator conditional distributions $q(W\mid R)$ is given by
@@ 4 Experiments
We confirmed the validity of our analytical results for inflated zero models by sampling data generating processes (DGPs), and numerical methods. In addition, we used our methods to perform sensitivity analyses on CLABSI data. Details of these experiments are in the Appendix. The code could be found at https://github.com/trungpq-ci/zero-inflation-bounds.
@@ 4.3 Data Application
Our working model is the proxy-augmented ZI MAR under assumptions A1 and A2, shown in Fig. 2 (b). Using the analytic bounds for the ZI MAR model derived in Section 3.4, we perform a sensitivity analysis to understand how the true CLABSI rate $p(X^{(1)}=1)$ changes as the proxy-indicator relationship $p(W\mid R)$ varies, within its compatibility range. First, we use the EM algorithm [Dempster et al., 1977] to maximize the observed data likelihood ^$\hat{\cal L}_{EM}(X,W,C)$ defined via the full data distribution consistent with our assumptions. Next, we invoke Theorem 5 to obtain the plug-in estimate for $p(W=0|R=1)$ and the bounds for $p(W=0|R=0)$. Finally, we do a grid search over the bounds interval, compute the full data distribution $p(R,X,W,C)$ for each value of $p(W=0|R=0)$ via (2) and obtain $p(X^{(1)})$ using standard g-formula adjustment in MAR models. The sensitivity analysis curve is shown in Fig 4. 222 This plot differs somewhat from the plot in the published version of the paper, due to a corrected data processing error. The conclusions on the underlying CLABSI rate were not substantially affected.
The values of $p(W=0|R=0)$ consistent with the model show that inability to make a CLABSI determination is strongly associated with access to patient data via electronic health records. For proxy EPIC, our obtained (sharp) bound for the nuisance parameter $p(W=0|R=0)$ is $[0.88,1]$, yielding the estimated range of the true CLABSI rate to be $[0.69,0.79]$. Compared with the baseline rate of $65\%$ under no-zero-inflation assumption, the rate’s bound implies that anywhere from $4\%$ to $14\%$ of true CLABSI cases are undercounted, even in our patient cohort with a highly elevated CLABSI prevalence.
We have repeated the analysis using the proxy-augmented ZI MAR model under assumptions A1 and A2, shown in Fig. 2 (c). In this case, bounds for $p(W=0|R=0,c)$ were obtained, for each value $c$. The narrowest bound $[0.985,1.0]$ corresponds to adult patients receiving outpatient parenteral antimicrobial therapy (OPAT) via a peripherally inserted central catheter (PICC). On the other hand, the widest bound $[0.03,1.0]$ corresponds to pediatric patients receiving chemotherapy via tunneled CVC, a type of catheter under the skin. We performed a search of $10^{6}$ points over the polytope comprised of these bounds and find the estimated range of true CLABSI rate to be $[0.77,0.93]$. That is, anywhere from 12%-28% of true CLABSI cases are undercounted.
All derived bounds for the true CLABSI rate were deemed to be medically plausible by our medical collaborators.
@@ Acknowledgements.
This research is funded in part by ONR N00014-21-1-2820, NSF 2040804, NSF CAREER 1942239, NIH R01 AI127271-01A1, AHRQ R01 HS027819.
@@ A.4.2 Bounds for ZI MCAR
Due to this lemma, an observed law $p(X,W)$ is consistent to the model if and only if $\forall x\neq 0,p(W=0,X=x)=p(W=0,X=1)$. We also require positivity $\forall x,p(X=x)>0$, so that $p(W\mid X)$ is well-defined.
Theorem 3 Consider a ZI MCAR model in Fig. 2 (a) (reproduced in Fig. 5) under proxy assumptions A1, A2, with categorical $X$ and binary $R,W$. Given a consistent observed law $p(X,W)$ satisfying positivity assumption, $\forall x,p(x)>0$, the set of compatible proxy-indicator conditionals $q(W\mid R)$ is given by
@@ A.5 ZI MAR proofs
Theorem 4 Consider a ZI MAR model in Fig. 2 (b) (reproduced in Fig. 6) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity assumption, $\forall x,c,p(x,c)>0$, the set of compatible proxy-indicator conditional distributions $q(W\mid R,C)$ is given by, for each value $c$,
@@
$q\left(X^{(1)},R,X,W,C\right)=q\left(X^{(1)},R,X,W\mid C\right)p(C).$ (43)
@@
Theorem 5 Consider a ZI MAR model in Fig. 2 (c) (reproduced in Fig. 7) under proxy assumptions A1 and A2, with categorical $X,C$ and binary $R,W$. Given a consistent observed law $p(X,W,C)$ satisfying positivity assumption, $\forall x,c,p(x,c)>0$, the set of compatible proxy-indicator conditional distributions $q(W\mid R)$ is given by